OpenStudy (anonymous):

Does my textbook have the wrong answer? The question is: given f:x--> e^2x and g:x--> 2x-1, find the defining equations of: a: ( f(inverse) operation g)(x) and b: (g operation f)^(inverse) (x). For q.A I got it right which was (1/2)In(2x-1); for q.B I got ((1/2)Inx+1)/2 but in my textbook the answer is (1/2)In((x+1)/2)

OpenStudy (anonymous):

gof is 2e^2x - 1 y = 2e^2x - 1 (y+1)/2 = e^2x [ln [(y+1)/2]]/2 = x interchange x and y

OpenStudy (anonymous):

oooh okay thankyou so much!

OpenStudy (anonymous):

click "good answer"

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you please help me to solve this problem? 3x/3-4/2